OpenStudy (anonymous):

True or false: sin x cos y = 1/2 (sin(x+y) + cos (x-y))

OpenStudy (anonymous):

its wrong.. so false

OpenStudy (anonymous):

Recall that \[\sin (x\pm y)=\sin x\cos y\pm \sin y\cos x\]